1. A business idea requires to spend $20,000 on training the staff this year and will save $1,000 a every year afterwards (because the staff is more productive).
a. Will the firm implement the idea if the interest rate is 4%? Explain.
b. Will the firm implement the idea if the interest rate is 8%? Explain.
c. At what interest rate is the firm indifferent between implementing the idea and abandoning it? Explain.
1
Expert's answer
2019-06-25T09:03:56-0400

Use NPV Cash outflow =20000

a) Cash inflow is the present value of perpetuity with annual payments of 1000 and r=4%

PV=C/r=1000/0.04=$25000

NPV=Cash inflow-Cash outflow=25000-20000=$5000

Implement the idea because of positive NPV

b) C=1000, r=8% Cash inflow =1000/0.08=$12500

NPV=12500-25000=-$12500

Do not implement; Negative NPV

c) 20000=1000/r

r=1000/20000=5%

at r=5% Cash inflow = cash outflow